OpenStudy (anonymous):

Integral of sqrt((1-cosx)/2) from [0,2pi] Can someone show me the steps to this?

OpenStudy (raden):

hint : sqrt((1-cosx)/2) = sqrt(sin^2 (x/2)) = sin(x/2)